Dabney Electronics currently has no debt. Its operating income (EBIT) is $30 million and its tax rate is 40 percent. It pays out all of its net income as dividends and has a zero growth rate. It has 2.5 million shares of stock outstanding. If it moves to a capital structure that has 40 percent debt and 60 percent equity (based on market values), its investment bankers believe its weighted average cost of capital would be 10 percent. What would its stock price be immediately after issuing debt if it changes to the new capital structure?

Hint: Find value of the firm after capitalization using Va = FCF1/(WACC-g), and then calculate price of the stock using P0 = [S + (D – D0) ] / n0)

a.

$90

b.

$60

c.

$48

d.

$200

e.

$72
Answer #1

Initially,

Weight of Debt = 0%

Weight of Equity = 100%

Now,

Weight of Debt = 40%

Weight of Equity = 60%

growth rate = 0%

tax rate = 40%

number of shares = 2.5 Mn

WACC = 10%

EBIT = 30 Mn

FCFE = EBIT(1-TAX) = 30(1-40%) = 18 Mn

Va = FCFE/(WACC-g) = 18/(10-0) = 180 Mn

Price = Value/Number of shares = 180 Mn/2.5 = 72

so, immediately after issuing debt of 40% the share price will be 72.

Ans: E) 72